Maggie is four times as old as her cousin Sam. In five years she will be six years less than three times his age. How old is Sam now?
Answer by FrankM(1040)

M=4S
M+5=3(S+5)-6
4S+5=3(S+5)-6
4S+5=3S+9
S=4 M=16